OpenStudy (anonymous):

how can you write 4 2/3 as an iproper fraction?

OpenStudy (anonymous):

14/3

OpenStudy (anonymous):

so how did u get that answer

OpenStudy (anonymous):

You know the full part is 3. And there are 4 full parts. So 3 * 4 = 12.

OpenStudy (anonymous):

sorry but i still dont get it

OpenStudy (anonymous):

In addition to your 4 'filled up' numbers you have 2/3 of a full number. So add 12 + 2.

OpenStudy (anonymous):

I simply must add that IMO there is nothing improper in 14/3 and this topic is just bullpellet.

OpenStudy (anonymous):

Try seeing a 1 as 3/3. And you have 4*1, so 4*3/3=12/3.

OpenStudy (anonymous):

Or use the cake metaphor : ) Imagine 4 cakes cut into 3 and 1 cake also cut into three, but with one piece eaten - hence 2/3 is still there.

OpenStudy (anonymous):